OpenStudy (anonymous):

Find the work W done by a force of 5 pound acting in the direction of 45 degree to the horizontal in moving an object 7 feet from (0,0) to (7,0) W = ? foot-pounds

OpenStudy (anonymous):

Also, this is dealing with vectors F and magnitude

OpenStudy (anonymous):

Ok, This is what I got. One sec

OpenStudy (mathmate):

\(work~done, W=\textbf{F.D}=FDcos(\theta)\) where \(\theta\) is the angle between \(\textbf{F}~and~\textbf{D}\). W is a scalar.

OpenStudy (mathmate):



OpenStudy (anonymous):

F = 5(cos ai + sin aj) cos 45 + sin 45 \[ \huge 5(\frac{\sqrt{2}}{2}i + \frac{\sqrt{2}}{2}j ) \] \[ \huge \frac{5\sqrt{2}}{2}i + \frac{5\sqrt{2}}{2}j \]

OpenStudy (anonymous):

Now I guess the W is the magnitude of force to the distance?

OpenStudy (anonymous):

Or magnitude time distance

OpenStudy (mathmate):

It is a dot product. So you need to put the distance as a vector as well, then proceed with the dot product!

OpenStudy (anonymous):

So..... \[\huge (\frac{5\sqrt{2}}{2}i + \frac{5\sqrt{2}}{2}j) * 7i \] \[\huge (\frac{5\sqrt{2}}{2}(7) + \frac{5\sqrt{2}}{2}(0)) \] \[\huge (\frac{5\sqrt{2}}{2}(7) + \frac{5\sqrt{2}}{2}(0)) \] \[\huge \frac{5\sqrt{2}}{2}(7) \] \[\huge W = \frac{35\sqrt{2}}{2} \] Correct?

OpenStudy (dan815):

yep that looks good

OpenStudy (anonymous):

Thanks!

OpenStudy (mathmate):

Yep, that's right.